# “Sharing, Not Showing Off”: How BeReal Approaches Authentic Self-Presentation on Social Media Through Its Design

### v1

@@ 4.1. Participants and Recruitment
We used a combination of convenience and snowball sampling to recruit participants. Initial recruitment efforts involved reaching out to any known BeReal users aged between 13 and 18, as well as contacting the university’s research participation pool to seek potential participants. Three participants were recruited from this effort. Recruiters were offered an additional $10 incentive for recruiting friends who met the study’s inclusion criteria. Three additional participants were recruited this way. To reach a broader audience, we posted ads on Instagram and Facebook targeting 13-18-year-olds residing in the United States. A total of 195 individuals responded and completed a screener survey. From this pool, 46 respondents met our inclusion criteria. We scheduled and conducted interviews with 29 of these respondents on a first-come, first-served basis.
@@
The mean age of the participants was 15.6 years (SD=1.54), with 22 (75.9%) identifying as girls, 6 (20.7%) as boys, and one (3.4%) as non-binary. The participants’ self-described racial/ethnic breakdown included 11 (37.9%) White/Caucasian, 9 (31%) Asian, 3 (10.3%) African American, 2 (6.9%) Hispanic, 1 (3.4%) Asian/White, 1 (3.4%) Hispanic/White, and 1 (3.4%) Asian/African. On average, participants reported having used BeReal for 5 months (mean=4.8 months, sd=1.7 months) at the time of the interview, excluding two novice users who had used BeReal for less than one month. To denote novice users with less than one month of experience with BeReal, add an “n” suffix to their participant name (e.g., “P10” as “P10n”). Approximately 58% of BeReal users in the U.S. identify as females (Curry, 2022). We do not have the racial or ethnic data for the overall population of BeReal users; our sample was predominantly White and Asian. The full list of participant information is available in Table 2.
@@
Table with information on the participant ID, age, gender, ethnicity, time since BeReal install, and number of friends on BeReal PID. Age Gender Ethnicity Time since BeReal install No. of friends on BeReal P01 14 Boy White/caucasian 3 months 30 P02 13 Girl White/caucasian 2 months 22 P03 14 Girl White/caucasian 3 months 20 P04 17 Girl White/caucasian 5 months 25-30 P05 13 Girl African American 2 months 20-25 P06 16 Girl White/caucasian 3 months 50+ P07 17 Girl White/caucasian 4 months 50+ (150)** P08n 17 Non-binary White/caucasian 1 week 1 P09 16 Girl Asian/Asian American 5 months 50+ P10n 18 Girl African American 1 week 8 P11 17 Boy White/caucasian 6 months 50+ (70-80) P12 17 Girl African American 4 months 50+ (70) P13 14 Boy White/caucasian 4 months 20-30 P14 16 Girl Asian/Asian American 4 months 20 P15 15 Girl Asian/Asian American 4 months 30 P16 13 Girl Asian/White 5 months 3 P17 15 Girl White/caucasian 6 months 30 P18 15 Girl Asian/Asian American 6 months 15 P19 17 Boy Hispanic 7 months 20-40 P20 18 Boy White/caucasian 7 months 50 P21 17 Girl Asian/Asian American 5 months 30-50 P22 15 Boy Hispanic 4 months 10-15 P23 16 Girl Asian/Asian American 6 months 50+ P24 16 Girl Hispanic/White 8 months 50 P25 17 Girl Asian/Asian American 8 months 70 P26 15 Girl Hispanic 5 months 50+ (100+) P27 13 Girl Asian/Asian American 6 months 50+ P28 15 Girl Asian/African 5 months 15 P29 17 Girl Asian/Asian American 2 months 10 * Novice users are denoted with (n). ** Numbers in brackets (N) for 50+ friends is the participants’ estimate of the number of friends they have.
@@ 4.4. Ethical Considerations and Researchers’ Positionality
The perspectives that participants shared with us suggest that our approach succeeds in yielding rich, authentic perspectives. Teens shared critical reflections on the norms of BeReal, admissions of judgment towards others, expressions of self-consciousness, and deviations from expected engagement patterns.
@@ 5.1. How BeReal’s Design Encourages Authentic Self-Presentation
For the most part, participants did not join BeReal because they were seeking a platform that promotes authentic self-presentation. Instead, participants said they joined for social reasons, saying things like, “all my friends were on it” (P03) or “my friend told me to” (P09). Others were intrigued by the platform after coming across it on Instagram or TikTok. However, participants shared that once they became BeReal users, the platform’s features influenced their self-presentation and led them to share frequently and authentically (specifically, to post unfiltered, unstaged photos of themselves). Participants identified four key ways that BeReal’s design promotes authentic self-presentation: discouraging staged posts, relieving posting pressure, encouraging interactions with smaller social circles, and maintaining a simple, mission-focused app design. We described each of these in more detail below.
@@
Participants explained that these features funnel them into very specific self-presentation choices. Some participants shared that they retake photos only when their “eyes are closed” or they “accidentally put [their] finger in the camera” (P11), not due to self-consciousness. They also avoided posting blank images, aiming to “always [have] something there” (P26), because blank posts “defeat the whole point” (P20). They said they try to post as soon as they receive a notification because waiting to post renders it “kind of pointless” (P04). They described the way BeReal encourages reciprocity and mutual sharing, saying, “You can’t see anyone else’s BeReal until you post yours for the day,” likening it to an “exchange” (P04). Given this context, posting a blank photo—withholding a piece of their lives—while being able to view others’ posts seemed like“cheating” (P13), “hypocritical” (P13), and “not fair” (P17) to several participants. Further, posting blank photos was seen as undermining or defeating the “point” (P15) or “purpose” of the app (P22). Participants said that when they see others post blank photos, their reactions range from mild disappointment to more pronounced disapproval. P13 said that a proliferation of blank photos would “ruin it” for them. Collectively, participants made clear that BeReal’s features communicate specific expectations for how they should engage with the platform. These designed expectations lead them to post unstaged photos and expect other users to do the same.
@@
Unlike most mainstream social media platforms, where sharing edited and curated content is the norm, BeReal’s designs allow users to share photos without the pressure to curate their posts by making sharing more low-stakes. For example, participants described the app-defined cadence of the daily posting by saying, “I think it does feel different just because I’m doing it so frequently, and because it has this connotation of, ‘okay, this doesn’t matter”’ (P24). Ephemerality further lowers the stakes for engaging with others. P16 shared: “I don’t feel as pressured to make it look super nice because it’s like it’ll just be away in 24 hours.” Moreover, the effort Realmojis demand and their ephemeral nature reduce the pressure to collect them. Unlike other platforms where “likes” signify social status, on BeReal, they do not carry much weight. As P15 put it, “There’s like no likes…I don’t think people usually really compare [the number of Realmojis they received].”
The effect of these features is that even compared to other casual online spaces like Snapchat, Instagram Close Friend Stories, and Finsta (Xiao et al., 2020), participants found BeReal to be more random, unlike other platforms where users put more effort into curating their content. P18 shared that they post photos on BeReal that they would not share on their Finsta because they are “just too boring.” P04 stated that Close Friends Stories were also considered less random than BeReal posts, as they are primarily used for sharing funny or noteworthy personal events. Because all users share whenever “the app tells [them] to do so” (P25), users feel comfortable under the shared vulnerability to post the “least exciting parts” (P27) of their day. P23 clarified the effect of the frequent, “low-fidelity” posting on BeReal stating, “[on other social media platforms] you have a purpose in posting these [posts]. BeReal, you’re not supposed to have a purpose in posting. You’re just posting just because of a random notification, throughout the whole day, would just tell you to post. There isn’t that intent behind it.” One participant made an analogy that BeReal “feels a lot like a Polaroid”, whereas Instagram “would be a picture hung on your walls—perfect and made to show off” (P29).
@@
Context collapse inevitably leads to curated self-presentation, and BeReal approached this phenomenon by encouraging users to curate their audience instead of their posts. Specifically, BeReal’s features encourage engagement with closer acquaintances and friends. This approach makes it easier for users to post casual, daily content without the fear of judgment, fostering an environment that promotes casual sharing rather than curated self-presentation. By enforcing reciprocity, the platform encourages users to limit their BeReal circles to those with whom they feel comfortable sharing their daily lives. As a result, users typically have fewer friends on BeReal than they do on other platforms. P04 explained that they add only the people they “talk to regularly or that someone that I would want to stay in touch with” because they wish to feel “comfortable with [friends] seeing parts of [their lives] that are personal.” Also, according to P10, scrolling through posts from less familiar friends on BeReal can become “overwhelming and just sort of boring.” Therefore, users tend to curate their friends, adding only those “that they care about seeing what they’re doing every single day” (P04). Moreover, by hiding friend counts, BeReal discourages popularity contests that can hinder authentic sharing. P04 explained: “On BeReal there’s definitely the least amount of pressure because nobody can see how many friends someone else has. Like, I don’t even know how many friends I have exactly…the number really doesn’t matter so you can have as few or as many as you want.” Overall, participants expressed that their friend lists on BeReal, in general, are “pretty much invite-only” (P22).
These features, which encourage audience curation, result in BeReal being described as“friendly” (P26),“cozy,” “home-like” (P16), and“communal” (P04). P18 noted,“Very few people will see my BeReal…it feels like there’s less pressure.” Participants also said that the BeReal experience made them“feel connected” (P28) to other users and appreciated the platform’s capacity to centralize close friends for updates that felt like“check-ins” (P14). Participants viewed the sense of a“judgment-free” (P16) environment that feels like a “safe, close space” (P29) fostered by a carefully curated audience as a defining characteristic that sets BeReal apart from other social media platforms.
@@
The simplicity of BeReal not only makes its purpose clear to users but also sets the platform apart from features like TikTok Now, which, despite its striking resemblance to BeReal, is perceived differently due to its integration with TikTok’s broader functionality. For example, P25 expressed that “I feel like most people know TikToks as those short 15-second videos, whereas BeReal’s its own thing. And so to merge [Now into TikTok] just feels a little bit more awkward and then more clunky, and then it just doesn’t feel as real as BeReal.” Also, P24 shared that the audience base in the apps is different; “TikTok has more functions, and I guess that you can’t really curate the friends that you interact with on TikTok as much as you can with BeReal.” TikTok Now is a part of TikTok, which to the participants has “always been a general video sharing app” and “it’s the same concept as BeReal but …they don’t talk as much about like being real, that’s not really their focus the same way that it is with the BeReal app” (P08n).
@@ 5.2.1. Validating an unfiltered life
The contrast between the parts of themselves that people share on BeReal and Instagram was particularly striking to participants. For example, P24 observed that BeReal is “breaking down the walls that Instagram has put up since it was launched. You’re seeing people at their best, and you’re seeing people edited. BeReal is trying to get people to understand that that’s not real.” This aligns with P06’s assessment, who shared that “all I see of them on Instagram is like traveling” and that they find it “refreshing” that on BeReal, they can see their friends “staying at home” (P06). P24 further explained: “If Instagram is a room of people at their best moment with heavy editing, BeReal is trying to get people at their worst moment with no chance of editing. BeReal is anti-Instagram, and Instagram is just people trying to show off their best selves.”
Participants viewed BeReal’s emphasis on an unfiltered life as not just beneficial in the immediate sense but as indicative of a broader potential shift in social media culture. For example, P25 described BeReal as “trying to bridge that gap between social media and real-life” by “bring[ing] more intimacy into the social media world,” and they enjoyed being a part of “a collective effort to like destigmatize like having unfiltered moments.” Further, P08 expressed the platform’s mission as “hopeful and inspiring …it was created to help teach people that like you don’t need to stress all the time about social media and being cool and like the way that people perceive you. It’s okay to be your real self.” P24 expressed appreciation for BeReal’s new take on social media by acknowledging how BeReal devalues overthinking and excessive value given to an Instagram post, stating “There’s so many people that overthink it and give just an Instagram post too much value and platforms like BeReal are it doesn’t matter. It’s just an app …That lightheartedness is really valuable in terms of social media.”
@@ 5.2.2. Motivating positive changes
A few participants also said that using BeReal encouraged them to be the person they aspire to portray online. For example, P02 stated: “I’ll feel guilty if I am not, like, not getting out of bed or if like, my BeReal comes along and I’m still in bed or something, but most of the time it’s, like, it’s nice like when I’m doing something productive, and then my BeReal comes along, and I’m like, ‘oh cool I get to be doing something interesting while I take my BeReal.”’ Rather than feeling overwhelmed by this, P02 stated that they feel “motivated” and that they want to be productive when their BeReal comes along. P24 echoed this perspective by saying that looking back through their BeReal posts served as “a little mirror” to them.
@@
Participants also observed that the norms and efforts toward authenticity on BeReal shielded them from the negative impacts of social comparison and approval anxiety often experienced on other platforms. P23 recognized the app’s focus on reducing the anxiety generated by metrics such as “likes,” “I don’t really get that many reactions either, but I don’t really care, so that’s the point of the app. I shouldn’t really care for what people are seeing or thinking about me when I’m posting it. That’s the point.” Furthermore, several participants noted that the changes they experienced on the app were self-directed, attributing this to the platform’s enjoyable and casual design, as highlighted by one participant who described it as “fun” (P06). For example, P04 shared, “I don’t think I’ve ever felt, like, forced …I feel like the whole idea is kind of, you’re not judging people, like, it’s not meant to be like that. It’s really just for fun, and you’re only on there if you want to be on there.”

### v2

@@ 4.1. Participants and Recruitment
We used a combination of convenience and snowball sampling to recruit participants. Initial recruitment efforts involved reaching out to any known BeReal users aged between 13 and 18, as well as contacting the university’s research participation pool to seek potential participants. Three participants were recruited from this effort. Participants were offered an additional $10 incentive for recruiting friends who met the study’s inclusion criteria. Three additional participants were recruited this way. To reach a broader audience, we posted ads on Instagram and Facebook targeting 13-18-year-olds residing in the United States. A total of 195 individuals responded and completed a screener survey. From this pool, 46 respondents met our inclusion criteria. We scheduled and conducted interviews with 29 of these respondents on a first-come, first-served basis.
@@
The mean age of the participants was 15.6 years (SD=1.54), with 22 (75.9%) identifying as girls, 6 (20.7%) as boys, and one (3.4%) as non-binary. The participants’ self-described racial/ethnic breakdown included 11 (37.9%) White/Caucasian, 9 (31%) Asian, 3 (10.3%) African American, 2 (6.9%) Hispanic, 1 (3.4%) Asian/White, 1 (3.4%) Hispanic/White, and 1 (3.4%) Asian/African. On average, participants reported having used BeReal for 5 months (mean=4.8 months, sd=1.7 months) at the time of the interview, excluding two novice users who had used BeReal for less than one month. To denote novice users with less than one month of experience with BeReal, add an “n” suffix to their participant name (e.g., “P10” as “P10n”). Approximately 58% of BeReal users in the U.S. identify as women (Curry, 2022). We do not have the racial or ethnic data for the overall population of BeReal users. The full list of participant information is available in Table 2.
@@
Table with information on the participant ID, age, gender, ethnicity, time since BeReal install, and number of friends on BeReal PID Age Gender Ethnicity Time since BeReal install No. of friends on BeReal P01 14 Boy White/Caucasian 3 months 30 P02 13 Girl White/Caucasian 2 months 22 P03 14 Girl White/Caucasian 3 months 20 P04 17 Girl White/Caucasian 5 months 25-30 P05 13 Girl African American 2 months 20-25 P06 16 Girl White/Caucasian 3 months 50+ P07 17 Girl White/Caucasian 4 months 50+ (150)** P08n 17 Non-binary White/Caucasian 1 week 1 P09 16 Girl Asian/Asian American 5 months 50+ P10n 18 Girl African American 1 week 8 P11 17 Boy White/Caucasian 6 months 50+ (70-80) P12 17 Girl African American 4 months 50+ (70) P13 14 Boy White/Caucasian 4 months 20-30 P14 16 Girl Asian/Asian American 4 months 20 P15 15 Girl Asian/Asian American 4 months 30 P16 13 Girl Asian/White 5 months 3 P17 15 Girl White/Caucasian 6 months 30 P18 15 Girl Asian/Asian American 6 months 15 P19 17 Boy Hispanic 7 months 20-40 P20 18 Boy White/Caucasian 7 months 50 P21 17 Girl Asian/Asian American 5 months 30-50 P22 15 Boy Hispanic 4 months 10-15 P23 16 Girl Asian/Asian American 6 months 50+ P24 16 Girl Hispanic/White 8 months 50 P25 17 Girl Asian/Asian American 8 months 70 P26 15 Girl Hispanic 5 months 50+ (100+) P27 13 Girl Asian/Asian American 6 months 50+ P28 15 Girl Asian/African 5 months 15 P29 17 Girl Asian/Asian American 2 months 10
Notes: Novice users are denoted with (n). Numbers in brackets (N) for 50+ friends is the participants’ estimate of the number of friends they have.
@@ 4.4. Ethical Considerations and Researchers’ Positionality
The perspectives that participants shared with us suggest that our approach succeeded in yielding rich, authentic perspectives. Teens shared critical reflections on the norms of BeReal, admissions of judgment towards others, expressions of self-consciousness, and deviations from expected engagement patterns.
@@ 5.1. How BeReal’s Design Encourages Authentic Self-Presentation
For the most part, participants did not join BeReal because they were seeking a platform that promotes authentic self-presentation. Instead, participants said they joined for social reasons, saying things like, “all my friends were on it” (P03) or “my friend told me to” (P09). Others were intrigued by the platform after coming across it on Instagram or TikTok. However, participants shared that once they became BeReal users, the platform’s features influenced their self-presentation and led them to share frequently and “authentically” (specifically, to post unfiltered, unstaged photos of themselves). Participants identified four key ways that BeReal’s design promotes authentic self-presentation: 1) discouraging staged posts, 2) relieving posting pressure, 3) encouraging interactions with smaller social circles, and 4) maintaining a simple, mission-focused app design. We describe each of these in more detail below.
@@
Participants explained that these features funnel them into very specific self-presentation choices. Some participants shared that they retake photos only when their “eyes are closed” or they “accidentally put [their] finger in the camera” (P11), not due to self-consciousness. They also said they avoid posting blank images, aiming to “always [have] something there” (P26), because blank posts “defeat the whole point” (P20). They said they try to post as soon as they receive a notification because waiting to post renders it “kind of pointless” (P04). They described the way BeReal encourages reciprocity and mutual sharing, saying, “You can’t see anyone else’s BeReal until you post yours for the day,” likening it to an “exchange” (P04). Given this context, posting a blank photo—withholding a piece of their lives—while being able to view others’ posts seemed like“cheating” (P13), “hypocritical” (P13), and “not fair” (P17) to some participants. Further, posting blank photos was seen as undermining or defeating the “point” (P15) or “purpose” of the app (P22). Participants said that when they see others post blank photos, their reactions range from mild disappointment to more pronounced disapproval. Collectively, participants made clear that BeReal’s features communicate specific expectations for how they should engage with the platform. These designed expectations lead them to post unstaged photos and expect other users to do the same.
@@
Unlike most mainstream social media platforms, where sharing edited and curated content is the norm, BeReal’s designs allow users to share photos without the pressure to curate their posts by making sharing more low-stakes. For example, participants described the app-defined cadence of the daily posting by saying, “I think it does feel different just because I’m doing it so frequently, and because it has this connotation of, ‘okay, this doesn’t matter”’ (P24). Ephemerality further lowers the stakes for engaging with others. P16 shared: “I don’t feel as pressured to make it look super nice because it’s like it’ll just be away in 24 hours.” Moreover, the effort Realmojis demand and their ephemeral nature reduce the pressure to collect them. Unlike other platforms where “Likes” signify social status, on BeReal, they do not carry much weight. As P15 put it, “There’s like no ‘Likes’…I don’t think people usually really compare [the number of Realmojis they received].”
The effect of these features is that even compared to other casual online spaces like Snapchat, Instagram Close Friend Stories, and Finsta (Xiao et al., 2020), participants found BeReal to be less crafted, unlike other platforms where users put more effort into curating their content. P18 shared that they post photos on BeReal that they would not share on their Finsta because they are “just too boring.” P04 stated that Close Friends Stories were also considered more curated than BeReal posts, as they are primarily used for sharing funny or noteworthy personal events. Because all users share whenever “the app tells [them] to do so” (P25), users feel comfortable posting the “least exciting parts” (P27) of their day. P23 clarified the effect of the frequent, low-fidelity posting on BeReal stating, “[on other social media platforms] you have a purpose in posting these [posts]. BeReal, you’re not supposed to have a purpose in posting. You’re just posting just because of a random notification, throughout the whole day, would just tell you to post. There isn’t that intent behind it.” One participant used the analogy that BeReal “feels a lot like a Polaroid,” whereas Instagram “would be a picture hung on your walls—perfect and made to show off” (P29).
@@
Context collapse inevitably leads to curated self-presentation, and BeReal attempted to address this phenomenon by encouraging users to curate their audience instead of their posts. Specifically, BeReal’s features encourage engagement with closer acquaintances and friends. This approach makes it easier for users to post casual, daily content without the fear of judgment, fostering an environment that promotes casual sharing rather than curated self-presentation. By enforcing reciprocity, the platform encourages users to limit their BeReal circles to those with whom they feel comfortable sharing their daily lives. As a result, users typically have fewer friends on BeReal than they do on other platforms. P04 explained that they add only the people they “talk to regularly or that someone that I would want to stay in touch with” because they wish to feel “comfortable with [friends] seeing parts of [their lives] that are personal.” Also, according to P10, scrolling through posts from less familiar friends on BeReal can become “overwhelming and just sort of boring.” Therefore, users tend to curate their friends, adding only those “that they care about seeing what they’re doing every single day” (P04). Moreover, by hiding friend counts, BeReal discourages popularity contests that can hinder authentic sharing. P04 explained: “On BeReal there’s definitely the least amount of pressure because nobody can see how many friends someone else has. Like, I don’t even know how many friends I have exactly…the number really doesn’t matter so you can have as few or as many as you want.” Overall, participants expressed that their friend lists on BeReal, in general, are “pretty much invite-only” (P22).
These features, which encourage audience curation, result in BeReal being described as “friendly” (P26), “cozy”, “home-like” (P16), and “communal” (P04). P18 noted, “very few people will see my BeReal…it feels like there’s less pressure.” Participants also said that the BeReal experience made them “feel connected” (P28) to other users and appreciated the platform’s focus on updates from close friends that felt like “check-ins” (P14). Participants described BeReal as a “judgment-free” (P16) environment that feels like a “safe, close space” (P29) fostered by a carefully curated audience as a defining characteristic that sets BeReal apart from other social media platforms.
@@
The simplicity of BeReal not only makes its purpose clear to users but also sets the platform apart from features like TikTok Now (TikTok, 2022), which, despite its striking resemblance to BeReal, is perceived differently because it is integrated into TikTok’s broader functionality. For example, P25 expressed that “I feel like most people know TikToks as those short 15-second videos, whereas BeReal’s its own thing. And so to merge [Now into TikTok] just feels a little bit more awkward and then more clunky, and then it just doesn’t feel as real as BeReal.” Also, P24 shared that the audience in each app is different; “TikTok has more functions, and I guess that you can’t really curate the friends that you interact with on TikTok as much as you can with BeReal.” TikTok Now is a part of TikTok, which to the participants has “always been a general video sharing app” and “it’s the same concept as BeReal but …they don’t talk as much about like being real, that’s not really their focus the same way that it is with the BeReal app” (P08n).
@@ 5.2.1. Validating an unfiltered life
The contrast between the parts of themselves that people share on BeReal versus Instagram was particularly striking to participants. For example, P24 observed that BeReal is “breaking down the walls that Instagram has put up since it was launched. You’re seeing people at their best, and you’re seeing people edited. BeReal is trying to get people to understand that that’s not real.” This aligns with P06’s assessment, who shared that “all I see of them on Instagram is like traveling” and that they find it “refreshing” that on BeReal, they can see their friends “staying at home” (P06). P24 further explained: “If Instagram is a room of people at their best moment with heavy editing, BeReal is trying to get people at their worst moment with no chance of editing. BeReal is anti-Instagram, and Instagram is just people trying to show off their best selves.”
Participants viewed BeReal’s emphasis on an unfiltered life as indicative of a broader potential shift in social media culture. For example, P25 described BeReal as “trying to bridge that gap between social media and real-life” by “bring[ing] more intimacy into the social media world,” and they enjoyed being a part of “a collective effort to like destigmatize like having unfiltered moments.” Further, P08 expressed the platform’s mission as “hopeful and inspiring …it was created to help teach people that like you don’t need to stress all the time about social media and being cool and like the way that people perceive you. It’s okay to be your real self.” P24 expressed appreciation for BeReal’s new take on social media by acknowledging how BeReal reduces the pressure, workload, and stress that come with posting on other platforms, stating “There’s so many people that overthink it and give just an Instagram post too much value and platforms like BeReal are it doesn’t matter. It’s just an app …That lightheartedness is really valuable in terms of social media.”
@@ 5.2.2. Motivating positive changes
A few participants also said that using BeReal encouraged them to be the person they aspire to portray online. For example, P02 stated: “I’ll feel guilty if I am not, like, not getting out of bed or if like, my BeReal notification comes along and I’m still in bed or something, but most of the time it’s, like, it’s nice like when I’m doing something productive, and then my BeReal comes along, and I’m like, ‘oh cool I get to be doing something interesting while I take my BeReal.”’ Rather than feeling overwhelmed by this, P02 stated that they feel “motivated” and that they want to be productive when their BeReal comes along. P24 echoed this perspective by saying that looking back through their BeReal posts served as “a little mirror” to them.
@@
Participants also observed that on BeReal they felt less anxiety about others’ perceptions and spent less time engaging in social comparison than they do on other platforms. P23 recognized the app’s focus on reducing the anxiety generated by metrics such as “Likes,” “I don’t really get that many reactions either, but I don’t really care, so that’s the point of the app. I shouldn’t really care for what people are seeing or thinking about me when I’m posting it. That’s the point.” Furthermore, several participants noted that the changes they experienced on the app were self-directed, attributing this to the platform’s enjoyable and casual design, as highlighted by one participant who described it as “fun” (P06). For example, P04 shared, “I don’t think I’ve ever felt, like, forced …I feel like the whole idea is kind of, you’re not judging people, like, it’s not meant to be like that. It’s really just for fun, and you’re only on there if you want to be on there.”